OpenStudy (anonymous):

I need serious help with these series of questions... My math teacher won't help me.

OpenStudy (anonymous):

http://i1341.photobucket.com/albums/o749/whalexnuker/untitled_zps80b97cb0.png I'm not looking for a direct answer (although I'll take one) so you can't report me or do whatever saying I'm going against the rules. I'm only asking for an explanation and help

terenzreignz (terenzreignz):

Some math teacher...

OpenStudy (anonymous):

@terenzreignz Ikr? :| My real math teacher is out of office and his sub won't help me either. All she said was "there's an attachment and that's what you submit" Well, okay, I already KNOW that. I don't understand the attachment.

terenzreignz (terenzreignz):

Okay, let's have a look at Splott's first, shall we?

terenzreignz (terenzreignz):

2x + 4 = -3x + 14

terenzreignz (terenzreignz):

Obviously, there can be nothing wrong with that^ step since it's what's given...

OpenStudy (anonymous):

Both of them start like that- I have to figure out which step from each is solved correctly and I don't know how to solve it... I tried going to wolframalpha and all that but no luck

OpenStudy (anonymous):

solved incorrectly* like, which steps are wrong

terenzreignz (terenzreignz):

Wolfie can't help you... all wolfie can do is give you the correct answer.

terenzreignz (terenzreignz):

For now, this is something only a human (and apparently, not a martian) can do. So let's get started ^_^

OpenStudy (anonymous):

No, that site can give me step by step

OpenStudy (anonymous):

thanks- I appreciate your time out

terenzreignz (terenzreignz):

If that site can give it to you step by step... what's the problem?

terenzreignz (terenzreignz):

Okay, anyway, let's look at the second step, which follows from the first... 2x + 4 + 3x = -3x + 14 + 3x anything wrong with this step?

OpenStudy (anonymous):

@terenzreignz I said the site /*CAN*/ give me step by step. The option is there but, obviously, i can't get step by step. I don't know. I said I didn't know how to solve it.

OpenStudy (dan815):

lol

terenzreignz (terenzreignz):

Okay... calm down. I asked you a question about that second step...

OpenStudy (anonymous):

I'm not upset or anything ._. I am calm. I know which step you asked the question on and, again, I don't know how to solve the original problem so how am I supposed to know

terenzreignz (terenzreignz):

From knowledge of the addition property of equality, I suppose. For any numbers a, b, and c, if a = b then a + c = b + c

OpenStudy (anonymous):

I'll be real with you, take me for a dummy. Math is NOT my best subject at all. It's very confusing. Speak each step to me, please. e_e

terenzreignz (terenzreignz):

Basically, you are allowed to add the same number to both sides of an equation. Which is exactly what occured in \[\Large 2x + 4 = -3x + 14\]\[\Large 2x +4\color{red}{+3x}= -3x + 14 \color{red}{+3x}\]

OpenStudy (anonymous):

Okay, that's in splott's but not fizzle's

terenzreignz (terenzreignz):

We're only looking at Splott's for now. So, since what Splott did in the second step is add the same number to both sides of the equation, then that step is valid, okay? ^_^

OpenStudy (anonymous):

Oh, okay. That's his first step aside from the original equation. I understand that step is correct

terenzreignz (terenzreignz):

Okay, the third step... \[\Large 2x + 4 + 3x = -3x+14 + 3x\]\[\Large \color{red}{5x}+4=14\]

terenzreignz (terenzreignz):

See what Splott did there?

OpenStudy (anonymous):

To the left side of the equal side, he/she/it added 3x+2x=5x and the 3x on the right side of the equal sign cancels out so the 14 is left

terenzreignz (terenzreignz):

Good. So, was it a legit step? ^_^

OpenStudy (anonymous):

The next step? I think it would be subtracting 4 from each side so 5x = 10?

terenzreignz (terenzreignz):

It should be... but I don't think Splott *subtracted* 4... do you? It looks more like he/she/it actually *added* 4 to both sides...

OpenStudy (anonymous):

yeah, he added it which messes up the rest of his work. from 5x=10... i get confused on having to either get rid of x or getting x by its self... is it -5 from each side?

terenzreignz (terenzreignz):

To be fair to Splott, adding 4 to both sides doesn't make the equation wrong... but it doesn't help in solving it neither. \[\Large 5x + 4 \color{red}{+4}= 14\color{red}{+4}\]

terenzreignz (terenzreignz):

The problem is the step AFTER that one... notice how the 4+4 on the left side became zero...

OpenStudy (dan815):